# Underwater Image Enhancement via Dehazing and Color Restoration

### v1

@@ 3.1.2 Channel Fusion Block (CFB)
Pixel-wise addition and concatenation are the most common strategies for multi-feature map fusion. However, these methods treat the information among different feature maps equally. We believe that due to the nonlinear absorption of light by water at different wavelengths, the information between channels in underwater image features is not equally important. Therefore, in our designed CFB, the information among different feature map channels is fused based on their importance. Fig. 3 shows the specific process of the CFB. For two given feature maps $x_{1}\in{{\mathbb{R}}^{c\times h\times w}}$ and $x_{2}\in{{\mathbb{R}}^{c\times h\times w}}$, we obtained the weights $\{{\alpha}_{1}$ and ${\alpha}_{2}\}$ using the global average pooling $GAP(\cdot)$, softmax function, and split operation. Then, we performed a weighted sum and refined the sum result using $PConv$, $DConv$, and residual connections. The more formal fusion process of the CFB is as follows:
@@ 3.2 Loss Function
In the experiment, to enhance the chromatic consistency of local regions, we set a sliding window with a size of $15\times 15$ and a stride of 1 to calculate the chromatic consistency loss for each local region. Finally, we obtained the total chromatic consistency loss by taking the mean of all local losses.
@@
Here, ${\lambda}_{1}$, ${\lambda}_{2}$, and ${\lambda}_{3}$ are constants that balance the proportion of each loss term, which were set as 3, 1, and 3, respectively, with numerous experiments.
@@ 4.1 Implementation Details
Our experiments were conducted on both synthetic underwater datasets [44] and real underwater dataset UIEBD [26]. In the synthetic underwater dataset, the reference images followed real color distributions, whereas the synthetic images simulated 10 different underwater degradation environments (I, IA, IB, II, and III simulated open sea underwater environments; 1, 3, 5, 7, and 9 simulated nearshore underwater environments). Each category of synthetic images had 1449 images, which were paired with reference images. We randomly selected 300 paired images from each category as the training set. The UIEBD dataset comprised 890 pairs of real underwater-reference images and 60 real underwater images without pairs. The reference images were obtained through enhancement using various traditional methods. We randomly partitioned 800 paired images as the training set. All images were resized to $256\times 256$ for training, and we augmented the dataset by applying random flips and rotations.
The experiments were conducted on tow NVIDIA GPUs (GeForce GTX 2080Ti) using the PyTorch 1.7.1 framework on an Ubuntu 16.04 system. We trained the network using the Adam optimizer with an initial learning rate of 0.001 and decayed it by a factor of $\frac{1}{2}$ every 50 epochs. We set the number of training iterations to 300 epochs and the batch size to 4.
@@ 4.2 Experiment settings
Benchmarks. In the synthetic underwater dataset, we randomly selected 30 additional paired images from each category of synthetic data as the test set, which was denoted as Test-S300. In the UIEBD dataset, we selected the remaining 90 paired images as the test set, which was denoted as Test-U90. Test-S300 and Test-U90 served as our full-reference benchmarks for the experiments. To demonstrate the generalization performance of our proposed network, we randomly selected 60 and 16 unpaired real underwater images from the UIEBD dataset and SQUID dataset [45], respectively, as the no-reference datasets. These datasets are called Test-U60 and Test-SQ16, respectively.
@@
Compared Methods. We compared the proposed WaterFormer to other state-of-the-art underwater image enhancement methods based on deep learning: UWCNN[44], Water-Net[26], Ucolor[28], FUnIE[29], U-shape[36], UDAformer[37], and LE-Net[46]. In addition, we compared WaterFormer with the following traditional methods: Fusion-based[19], Retinex-based[21], MLLE[14], UDCP[47], RED[25], and HFM[48].
@@
Evaluation Metrics. For Test-S300 and Test-U90, we used SSIM, PSNR, and NRMSE as the full-reference evaluation metrics for our experiments. These metrics are widely used to assess the similarity between two images due to their convenience and accuracy. Higher values of SSIM and PSNR or lower values of NRMSE indicate greater similarity between two images. For Test-U60, we used UCIQEE [49] and UIQM [50] as the no-reference evaluation metrics. Higher values of UCIQE and UIQM indicate better visual perception of the underwater images by humans. However, UCIQE and UIQM may not accurately reflect the actual visual perception in certain cases.
Test-S300 Test-U90 Overhead
Methods SSIM↑ PSNR↑ NRMSE↓ SSIM↑ PSNR↑ NRMSE↓ #Param↓ MACs↓ Time↓
Fusion-based[19] 0.651 12.65 0.685 0.872 21.80 0.188 × × 0.073s
Retinex-based[21] 0.641 14.54 0.520 0.773 17.34 0.283 × × 0.208s
MLLE[14] 0.603 12.38 0.667 0.740 16.50 0.315 × × 0.038s
UDCP[47] 0.597 13.84 0.500 0.689 13.85 0.418 × × 1.694s
RED[25] 0.659 12.66 0.674 0.833 19.03 0.244 × × 0.020s
HFM[48] 0.676 15.00 0.377 0.835 18.32 0.293 × × 0.328s
UWCNN[44] 0.640 14.62 0.456 0.818 18.10 0.265 0.041M 2.649G 0.002s
Water-Net[26] 0.746 19.25 0.322 0.905 23.17 0.153 1.091M 71.42G 0.016s
FUnIE[29] 0.769 21.34 0.231 0.814 18.94 0.239 7.020M 10.24G 0.002s
Ucolor[28] 0.813 22.23 0.235 0.897 21.23 0.180 148.8M 404.8G 0.169s
U-shape[36] 0.793 21.89 0.213 0.806 20.28 0.208 22.82M 2.983G 0.025s
UDAformer[37] 0.798 22.39 0.226 0.921 23.38 0.147 9.594M 41.59G 0.042s
LE-Net[46] 0.774 19.19 0.277 0.865 21.42 0.180 0.014M 0.690G 0.006s
WaterFormer 0.907 29.40 0.116 0.929 24.39 0.137 0.313M 7.749G 0.023s
@@
Table 1: Quantitative comparison results for each method on the Test-S300 and Test-U90 datasets. Red and blue colors indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
4.3 Quantitative Comparisons
Full-reference Evaluation. Tab. 1 shows the quantitative comparison results on Test-S300 and Test-U90. The optimal results are highlighted in red, and the second-best results are highlighted in blue. WaterFormer achieved the best SSIM, PSNR, and NRMSE metrics on both full-reference test sets. Notably, for Test-S300, WaterFormer surpassed the second-best method by 9.4% in SSIM, 31.6% in PSNR, and shows a 9.7% lower NRMSE. This significant improvement indicated that previous methods were not able to effectively restore the true color distribution, detail distribution, and noise distribution of underwater images. Similarly, for Test-U90, where different methods were used to enhance the original images to obtain reference images, WaterFormer surpassed the second-best method by 0.8% in SSIM, 4.3% in PSNR, and exhibited a 1.0% lower NRMSE. This represents that our proposed network performs better in fitting underwater image enhancement compared with the previously best-performing methods.
Overhead. In Tab. 1, we included the inference speed of each method. Additionally, for deep learning methods, we measured the overall cost of the network using the total number of parameters (#Param) and multiply-accumulate operations (MACs). All measurements of the inference speed and network cost were conducted on images of size $256\times 256$. Compared to other methods, our proposed WaterFormer achieves optimal inference results while maintaining fast inference speed and relatively low network overhead.
@@ 4.3 Quantitative Comparisons
Non-reference Evaluation. Tab. 2 shows the quantitative comparison results on Test-U60 and Test-SQ16. The best-performing and second-best results are highlighted in red and blue, respectively. Unfortunately, our proposed WaterFormer did not achieve the best or second-best results in the UIQM and UCIQE metrics on either no-reference datasets. As referenced  [28, 36] notes, while the UIQM and UCIQE metrics effectively evaluate the color, brightness, contrast, saturation, and sharpness of underwater images, they struggle to accurately measure color artifacts and shifts. In scenarios with more color artifacts and shifts, these metrics tend to yield higher scores, contradicting human perception. Our WaterFormer consistently avoids color artifacts and shifts, contributing to our lower UIQM and UCIQE scores. Hence, while we consider UIQM and UCIQE scores as indicative, they shouldn’t be the sole criteria for assessing the algorithm performance.
Test-U60 Test-SQ16
Methond UIQM↑ UCIQE↑ UIQM↑ UCIQE↑
@@
LE-Net[46] 2.549 0.561 2.297 0.512
WaterFormer 2.546 0.594 2.372 0.543
Table 2: Quantitative comparison results for each method on the Test-U60 and Test-SQ16 datasets. Red and blue colors indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
4.4 Qualitative Comparisons
Full-reference Evaluation. Fig. 4 and Fig. 5 show the qualitative comparison results on Test-S300 and Test-U90, respectively. For the Test-S300 dataset, the images enhanced by Fusion-based and RED methods had a reddish bias and failed to recover the true image details from the haze. The Retinex-based and UWCNN methods reduced the saturation of the enhanced images and even distorted the true color tones. Water-Net and Ucolor approximated the real colors of the synthesized images, but they still introduced significant artifacts. In contrast, WaterFormer did not generate artifacts and closely restored the brightness, color tones, and details in each region of the recovered images. Regarding the Test-U90 dataset, apart from Water-Net and WaterFormer, other methods failed to recover the desired colors, introduced artifacts, or could not effectively dehaze the images. Our proposed WaterFormer could better restore the colors and details and produced more accurate results than Water-Net.
Non-reference Evaluation. Fig. 6 and Fig. 7 show the qualitative comparison results on Test-U60 and Test-SQ16 , respectively. For the Test-U60 dataset, none of the methods satisfactorily performed. For example, the Fusion-based method generated artificial colors, the UDCP method produced low-brightness results, the UWCNN method shows minimal enhancement effect, the Water-Net method had limited effect on the shark image in the third row, the FUnIE method tended to produce color artifacts, and the U-shape method yielded low contrast results. Meanwhile, the images enhanced by WaterFormer achieved satisfying brightness, contrast, and color. Similarly, for the Test-SQ16 dataset, the Fusion-based, UDCP, UWCNN, and FUnIE methods failed to effectively restore the image colors. Water-Net and FUnIE introduced color shifts and artifacts, whereas the U-shape method had disappointing contrast and edge details in the enhanced images. In contrast, the images enhanced by WaterFormer restored as many colors and details as possible but exhibited minimal color shifts and artifacts.
@@ 4.4 Qualitative Comparisons
Fig. 4: Qualitative Comparison Results on Test-S300. (a) Input, (b) Fusion-based[19], (c) Retinex-based[21], (d) RED[25], (e) UWCNN[44], (f) Water-Net[26], (g) Ucolor[28], (h) WaterFormer, (i) GT.
@@
Fig. 5: Qualitative Comparison Results on Test-U90. (a) Input, (b) Fusion-based[19], (c) Retinex-based[21], (d) RED[25], (e) UWCNN[44], (f) Water-Net[26], (g) Ucolor[28], (h) WaterFormer, (i) GT.
@@
Fig. 6: Qualitative Comparison Results on Test-U60. (a) Input, (b) Fusion-based[19], (c) UDCP[47], (d) UWCNN[44], (e) Water-Net[26], (f) FUnIE[29], (g) U-shape[36], (h) WaterFormer.
@@
Fig. 7: Qualitative Comparison Results on Test-SQ16. (a) Input, (b) Fusion-based[19], (c) UDCP[47], (d) UWCNN[44], (e) Water-Net[26], (f) FUnIE[29], (g) U-shape[36], (h) WaterFormer.
[…]
Fig. 8: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-S300, respectively. (a) Input, (b) MLLE[14], (c) HFM[48], (d) Ucolor[28], (e) UDAformer[37], (f) LE-Net[46], (g)WaterFormer, (h) GT.
[…]
Fig. 9: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-U90, respectively. (a) Input, (b) MLLE[14], (c) HFM[48], (d) Ucolor[28], (e) UDAformer[37], (f) LE-Net[46], (g)WaterFormer, (h) GT.

### v2

@@ 3.1.2 Channel Fusion Block (CFB)
Pixel-wise addition and concatenation are the most common strategies for multi-feature map fusion. However, these methods treat information among different feature maps equally. We believe that due to the nonlinear absorption of light by water at different wavelengths, the information between channels in underwater image features is not equally important. Therefore, in our designed CFB, the information among different feature map channels is fused based on their importance. Fig. 3 shows the specific process of the CFB. For two given feature maps $x_{1}\in{{\mathbb{R}}^{c\times h\times w}}$ and $x_{2}\in{{\mathbb{R}}^{c\times h\times w}}$, we obtained the weights $\{{\alpha}_{1}$ and ${\alpha}_{2}\}$ using the global average pooling $GAP(\cdot)$, softmax function, and split operation. Then, we performed a weighted sum and refined the sum result using $PConv$, $DConv$, and residual connections. The more formal fusion process of the CFB is as follows:
@@ 3.2 Loss Function
In the experiment, to enhance the chromatic consistency of local regions, we set a sliding window with a size of $15\times 15$ (see Tab. 7) and a stride of 1 to calculate the chromatic consistency loss for each local region. Finally, we obtained the total chromatic consistency loss by taking the mean of all local losses.
@@
Here, ${\lambda}_{1}$, ${\lambda}_{2}$, and ${\lambda}_{3}$ are constants that balance the proportion of each loss term, set to 1, 1, and 2, respectively, based on numerous experiments (see Tab. 8).
@@ 4.1 Implementation Details
Our experiments were conducted on both synthetic underwater datasets [12] and real underwater dataset UIEBD [13]. In the synthetic underwater dataset, the reference images followed real color distributions, whereas the synthetic images simulated 10 different underwater degradation environments (I, IA, IB, II, and III simulated open sea underwater environments; 1, 3, 5, 7, and 9 simulated nearshore underwater environments). Each category of synthetic images had 1449 images, which were paired with reference images. We randomly selected 300 paired images from each category as the training set. The UIEBD dataset includes 890 paired underwater real images and reference images, as well as 60 unpaired underwater real images. The reference images were obtained through enhancement using various traditional methods. We randomly partitioned 800 paired images as the training set. All images were resized to $256\times 256$ for training, and we augmented the dataset by applying random flips and rotations.
The experiments were conducted on tow NVIDIA GPUs (GeForce GTX 2080Ti) using the PyTorch 1.7.1 framework on an Ubuntu 16.04 system. We adopt the AdamW optimizer with hyperparameters ${\beta}_{1}=0.9$ and ${\beta}_{2}=0.999$, initializing the learning rate at $1\times 10^{-3}$. A cosine annealing scheduler (CosineAnnealingLR) progressively adjusts the learning rate from $1\times 10^{-3}$ to $1\times 10^{-7}$ over 400 training epochs, with a batch size of 4.
@@ 4.2 Experiment settings
Benchmarks. In the synthetic underwater dataset, we randomly selected 30 additional paired images from each synthetic data category as the test set, which was denoted Test-S300. In the UIEBD dataset, we selected the remaining 90 paired images as the test set, which was denoted as Test-U90. Test-S300 and Test-U90 served as our full-reference benchmarks for the experiments. To demonstrate the generalization performance of our proposed network, we randomly selected 60 and 16 unpaired real underwater images from the UIEBD dataset and SQUID dataset [37], respectively, as the no-reference datasets. These datasets are called Test-U60 and Test-SQ16, respectively.
@@
Compared Methods. We compared the proposed WaterFormer to other state-of-the-art UIE methods based on deep learning: UWCNN[12], Water-Net[13], Ucolor[23], FUnIE[25], U-shape[30], UDAformer[31], GuidedHybSensUIR[38], DM-water[27], and WF-Diff[28]. In addition, we compared WaterFormer with the following traditional methods: Fusion-based[18], Retinex-based[19], MLLE[15], UDCP[39], RED[21], and HFM[40].
@@
Evaluation Metrics. For Test-S300 and Test-U90, we used SSIM, PSNR, and NRMSE as the full-reference evaluation metrics for our experiments. These metrics are widely used to assess the similarity between two images due to their convenience and accuracy. Higher values of SSIM and PSNR or lower values of NRMSE indicate greater similarity between two images. For Test-U60, we used UCIQEE [41] and UIQM [42] as no-reference evaluation metrics. Higher values of UCIQE and UIQM indicate better visual perception of the underwater images by humans. However, UCIQE and UIQM may not accurately reflect actual visual perception in certain cases.
Methods Test-S300 Test-U90 Overhead
SSIM↑ PSNR↑ NRMSE↓ SSIM↑ PSNR↑ NRMSE↓ #Param(M)↓ MACs(G)↓ Time(s)↓
Fusion-based[18] 0.651 12.65 0.685 0.872 21.80 0.188 × × 0.073
Retinex-based[19] 0.641 14.54 0.520 0.773 17.34 0.283 × × 0.208
MLLE[15] 0.603 12.38 0.667 0.740 16.50 0.315 × × 0.038
UDCP[39] 0.597 13.84 0.500 0.689 13.85 0.418 × × 1.694
RED[21] 0.659 12.66 0.674 0.833 19.03 0.244 × × 0.020
HFM[40] 0.676 15.00 0.377 0.835 18.32 0.293 × × 0.328
UWCNN[12] 0.640 14.62 0.456 0.818 18.10 0.265 0.041 2.649 0.002
Water-Net[13] 0.746 19.25 0.322 0.905 23.17 0.153 1.091 71.42 0.016
FUnIE[25] 0.769 21.34 0.231 0.814 18.94 0.239 7.020 10.24 0.002
Ucolor[23] 0.813 22.23 0.235 0.897 21.23 0.180 148.8 404.8 0.169
[…]
UDAformer[31] 0.798 22.39 0.226 0.921 23.38 0.147 9.594 41.59 0.042
[…]
WF-Diff[28] 0.827 22.41 0.234 0.924 23.82 0.144 18.48 122.76 0.465
[…]
@@
Table 2: Quantitative comparison results for each method on the Test-S300 and Test-U90 datasets. Bold and underline indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
4.3 Results and Analysis
Full-reference Evaluation. Tab. 2 shows the quantitative comparison results on Test-S300 and Test-U90. The optimal results are presented in bold, while the second-best results are marked with an underline. WaterFormer achieved excellent SSIM, PSNR, and NRMSE scores on both full-reference test sets. What is noteworthy is that, for the Test-S300 dataset, WaterFormer exceeded the second-best method by 1.66% in SSIM, 13.3% in PSNR, and achieved a 19.8% lower NRMSE. This significant improvement indicated that previous methods were not able to effectively restore the true color distribution, detail distribution, and noise distribution of underwater images. For the Test-U90 dataset, WaterFormer achieved the best performance in SSIM, surpassing the second-best method, DM-water, by 0.22%. However, in terms of PSNR and NRMSE, WaterFormer ranked second, being 2.65% lower and 0.76% higher than DM-water, respectively. SSIM considers the consistency of brightness, contrast, and structural information, indicating that WaterFormer focused more on results aligned with human visual perception. In contrast, PSNR and NRMSE are more sensitive to pixel-level noise, suggesting a greater emphasis on noise reduction capability. Despite slight differences in their inferred results, the evaluation metrics were very close, demonstrating WaterFormer’s exceptional ability in image enhancement and fitting.
[…]
Fig. 4: Qualitative Comparison Results on Test-S300. (a) Input, (b) Fusion-based[18], (c) Retinex-based[19], (d) Water-Net[13], (e) Ucolor[23], (f) WF-Diff[28], (g) WaterFormer, (h) GT.
[…]
Fig. 5: Qualitative Comparison Results on Test-U90. (a) Input, (b) Fusion-based[18], (c) Retinex-based[19], (d) Water-Net[13], (e) Ucolor[23], (f) WF-Diff[28], (g) WaterFormer, (h) GT.
Fig. 4 and Fig. 5 show the qualitative comparison results on Test-S300 and Test-U90, respectively. For the Test-S300 dataset, images enhanced using Fusion-based and Retinex-based methods exhibited noticeable reddish and blackish color biases, failing to recover the true details obscured by haze. Although the Water-Net, Ucolor, and WF-Diff methods succeeded in roughly restoring the original colors of the synthesized images, they still fell short of recovering the full saturation and tonal distribution. Furthermore, they introduced color artifacts that were absent in the original images. In contrast, WaterFormer did not produce such artifacts and successfully restored brightness, color tones, and details across different regions of the enhanced images. Regarding the Test-U90 dataset, Fusion-based and Retinex-based enhancements still resulted in red and black color biases, along with prominent red edge artifacts. The ability of Ucolor to recover colors was relatively limited, and the enhanced images lacked sufficient tone and saturation. However, the results from Water-Net, WF-Diff, and WaterFormer were closer to the reference images, demonstrating their effectiveness in accurately restoring both color and fine details from real underwater scenes.
@@ 4.3 Results and Analysis
Non-reference Evaluation. Tab. 3 shows the quantitative comparison results on Test-U60 and Test-SQ16. The best and second best results are presented in bold and underlined, respectively. Unfortunately, our proposed WaterFormer did not achieve the best or second-best results in the UIQM and UCIQE metrics on either no-reference datasets. As noted in  [23, 30], while the UIQM and UCIQE metrics effectively evaluate the color, brightness, contrast, saturation, and sharpness of underwater images, they struggle to accurately measure color artifacts and shifts. In scenarios with more color artifacts and shifts, these metrics tend to yield higher scores, contradicting human perception. Our WaterFormer consistently avoids color artifacts and shifts, contributing to our lower UIQM and UCIQE scores. Hence, while we consider the UIQM and UCIQE scores as indicative, they should not be the sole criteria for assessing the algorithm performance.
Methods Test-U60 Test-SQ16
UIQM↑ UCIQE↑ UIQM↑ UCIQE↑
@@
DM-water[27] 2.704 0.575 2.212 0.525
WF-Diff[28] 2.785 0.576 2.336 0.520
GuidedHybSensUIR[38] 2.757 0.584 2.440 0.556
WaterFormer 2.594 0.581 2.395 0.551
Table 3: Quantitative comparison results for each method on the Test-U60 and Test-SQ16 datasets. Bold and underline indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
@@
Fig. 6: Qualitative Comparison Results on Test-U60. (a) Input, (b) UDCP[39], (c) MLLE[15], (d) FUnIE[25], (e) U-shape[30], (f) DM-water[27], (g) GuidedHybSensUIR[38], (h) WaterFormer.
@@
Fig. 7: Qualitative Comparison Results on Test-SQ16. (a) Input, (b) UDCP[39], (c) MLLE[15], (d) FUnIE[25], (e) U-shape[30], (f) DM-water[27], (g) GuidedHybSensUIR[38], (h) WaterFormer.
Fig. 6 and Fig. 7 show the qualitative comparison results on Test-U60 and Test-SQ16, respectively. For the Test-U60 dataset, none of the methods produced satisfactory results. For example, while the UDCP method performed color correction, its capability was relatively weak, and the resulting images suffered from insufficient brightness. The MLLE method, on the other hand, produced incorrect color recovery results, such as a green submarine body (second row) and a purple ocean (third row). The FUnIE method introduced color artifacts that were inaccurate. All of the U-shape, DM-water, and GuidedHybSensUIR methods produced enhanced images that lacked adequate contrast and sharpness. Unlike the above methods, the images restored by WaterFormer exhibited satisfactory brightness, contrast, sharpness, and color. Similarly, for the Test-SQ16 dataset, the UDCP method again produced images with insufficient brightness. The MLLE and FUnIE methods struggled with recovering the correct colors and introduced severe color artifacts. The U-shape, DM-water and GuidedHybSensUIR methods also failed to restore the contrast and sharpness of the images adequately. In comparison, the images generated by WaterFormer restored more colors and details as much as possible, without introducing any additional color shifts or artifacts.
Overhead. In Tab. 2, we included the inference speed of each method. Furthermore, for deep learning methods, we measured the total cost of the network using the total number of parameters (#Param) and multiply-accumulate operations (MACs). All inference speed and network cost measurements were conducted on images of size $256\times 256$. Compared to other methods, our proposed WaterFormer achieves optimal inference results while maintaining fast inference speed and relatively low network overhead.
@@
Fig. 8: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-S300, respectively. (a) Input, (b) MLLE[15], (c) HFM[40], (d) UDAformer[31], (e) DM-water[27], (f) GuidedHybSensUIR[38], (g)WaterFormer, (h) GT.
@@
Fig. 9: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-U90, respectively. (a) Input, (b) MLLE[15], (c) HFM[40], (d) UDAformer[31], (e) DM-water[27], (f) GuidedHybSensUIR[38], (g)WaterFormer, (h) GT.